OpenStudy (cypress7):

Please help I will medal to the first person who answers what is 4 1/5 + 3 1/10 A. 7 2 /15 B. 7 3/10 C 7 1 /50 D.7 1/5

OpenStudy (cypress7):

umm lets see 4 1/5 would equal too 21/5 I think and 3 1/10 would be 31/10 I think

OpenStudy (anonymous):

First write these two fractions in standart form

OpenStudy (anonymous):

4 1/5=20/5+1/5=21/5

OpenStudy (anonymous):

3 1/10=30/10+1/10=31/10

OpenStudy (anonymous):

Now write fraction 21/5 as (21*2)/(5*2)=42/10

OpenStudy (anonymous):

Now you can add these two fractions : 42/10+31/10=73/10

OpenStudy (anonymous):

write 73/10 in mixed form: 70/10+3/10=7+3/10=7 3/10

OpenStudy (cypress7):

So the correct answer is B right?

OpenStudy (anonymous):

So the answer is B

OpenStudy (cypress7):

Thank you! I appreticate that you help me go through the question